# Partial Regularity for the three-dimensional Stochastic Ericksen–Leslie equations

### v1

@@ 3. Martingale suitable weak solution
Let $T>0$ and $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies (3.6). Then we say $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (3.3) in the sense of distribution on $Q_{T}:=\mathbb{T}^{3}\times(0,T]$ subject to ${{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}$, and for any $\phi_{1}\in C_{0}^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, $\mathop{\rm div}\nolimits\phi_{1}=0$, $\phi_{2}\in C_{0}^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, and $t>0$, the following weak formulation of (3.3) is satisfied:
@@
Now we are ready to introduce the definition of suitable weak solutions:
Definition 3.2.
Let $T>0$ and $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies (3.6). Then a suitable weak solution to (1.1) is a triple $({{\bf{u}}},P,{{\bf{d}}})$ such that if ${{\bf{v}}}={{\bf{u}}}-{{\bf{z}}}$ and ${\pi}=P-{Q}$ where $({{\bf{z}}},{Q})$ is the solution to (3.2), then
  1. (1)
@@
  3. (3)
    $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (3.3) in the sense of distribution on $Q_{T}$ as specified in Def. 3.1,
  4. (4)
    for all $t\leq T$ and almost every $s<t$,
    $\displaystyle\int_{\mathbb{T}^{3}\times\{t\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)+\int_{s}^{t}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}$
    $\displaystyle\leq\int_{\mathbb{T}^{3}\times\{s\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)$
    $\displaystyle+\int_{s}^{t}\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]+({{\bf{z}}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
  5. (5)
    for any $\varphi\in C_{0}^{\infty}(Q_{T})$, $\varphi\geq 0$,
    $\displaystyle\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{Q_{T}}(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2})\varphi$
    $\displaystyle\leq\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{\bf{v}}]\cdot{\bf{z}}\varphi$
    $\displaystyle+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{\bf{v}}\cdot{\bf{z}}\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
    $\displaystyle+\int_{Q_{T}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{\bf{d}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
    $\displaystyle+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla{{\bf{d}}})\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
    $\displaystyle-\int_{Q_{T}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{Q_{T}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Next, we introduce the notions of martingale suitable weak solutions.
Definition 3.3.
A martingale suitable weak solution to (1.1) is a quadruple $(\widetilde{\mathfrak{B}},\widetilde{{\bf{u}}},\widetilde{P},\widetilde{{\bf{d}}})$, where $\widetilde{\mathfrak{B}}$ is a stochastic basis
(3.7) $\widetilde{\mathfrak{B}}=(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t))_{t\geq 0}),$
and $(\widetilde{W}(t))_{t\geq 0}$ is a Brownian motion adapted to the filtration {~$\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with values in $D(A^{\delta})$ and covariance operator $\mathcal{L}$, such that
@@
is an ~$\widetilde{\mathcal{F}}_{t}$ adapted process, and
@@
is a measurable mapping, and there exists a set $\widetilde{\Omega}_{0}\subset\widetilde{\Omega}$ of full probability such that on this set the triple $(\widetilde{{\bf{u}}}(\cdot,\widetilde{\omega}),\widetilde{P}(\cdot,\widetilde{\omega}),\widetilde{{\bf{d}}}(\cdot,\widetilde{\omega}))$ is a suitable weak solution in the sense of Definition 3.2.
To state the main result of our paper, we introduce the notion of singular points for a suitable weak solution.
Definition 3.4.
@@
Theorem 3.3.
With the assumptions (3.1), suppose furthermore that the Brownian motion $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\frac{1}{4}+\delta})$ for some $\delta>0$. Then there exists a martingale suitable weak solution to (1.1) such that with full probability,
@@
4. Existence
@@
In this section, we will establish the existence of martingale suitable weak solutions to (1.1) in the sense of Definition 3.3.
4.1. Formal derivation of energy inequalities
The following a priori energy estimate will play an essential role in subsequent analysis.
Lemma 4.1 (Local energy equality).
Fix $\omega\in\Omega$ with ${{\bf{z}}}(\omega)\in L^{\infty}(0,T;H)\cap L^{2}(0,T;V)$ and let the triple $({{\bf{v}}}(\omega),{\pi}(\omega),{{\bf{d}}}(\omega))$ be a classical solution to (3.3), then for any $\varphi\in C_{0}^{\infty}(Q_{T})$, $\varphi\geq 0$, it holds that
(4.1) $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\varphi+\int_{\mathbb{T}^{3}}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2}\right)\varphi$
$=\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{{\bf{v}}}}\cdot{{{\bf{z}}}}\varphi$
$+\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{{\bf{v}}}\cdot{{\bf{z}}}\right)({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
$+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$+\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
$-\int_{\mathbb{T}^{3}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi$
Proof.
Taking the inner product of (3.3)${}_{1}$ with respect to ${{\bf{v}}}\varphi$ and integrating over $\mathbb{T}^{3}$ gives
$\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{v}}}\cdot{{\bf{v}}}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\partial_{t}\varphi.$
Using integration by parts we can derive
$\displaystyle\int_{\mathbb{T}^{3}}-\Delta{{\bf{v}}}\cdot{{\bf{v}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}:\nabla({{\bf{v}}}\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla{{\bf{v}}}:({{\bf{v}}}\otimes\nabla\varphi)$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi+\int_{\mathbb{T}^{3}}\nabla\left(\frac{|{{\bf{v}}}|^{2}}{2}\right)\cdot\nabla\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}\Delta\varphi.$
Now, we can see
$\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla({{\bf{z}}}+{{\bf{v}}})]\cdot{{\bf{v}}}\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{z}}}]\cdot{{\bf{v}}}\varphi+\int_{\mathbb{T}^{3}}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\left(\frac{|{{\bf{v}}}|^{2}}{2}\right)\varphi$
$\displaystyle=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{z}}}]\cdot{{\bf{v}}}\varphi-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]({\bf{v}}\cdot{\bf{z}})-\int_{\mathbb{T}^{3}}\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]({\bf{v}}\cdot{\bf{z}})+\frac{|{{\bf{v}}}|^{2}}{2}({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]\cdot{{\bf{z}}}\varphi-\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+({\bf{v}}\cdot{\bf{z}})\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi].$
Meanwhile, we have
$\int_{\mathbb{T}^{3}}\nabla{\pi}\cdot{{\bf{v}}}\varphi=-\int_{\mathbb{T}^{3}}{\pi}{{\bf{v}}}\cdot\nabla\varphi.$
For the Ericksen stress tensor term, using (3.4) we obtain
$\displaystyle\int_{\mathbb{T}^{3}}-\mathop{\rm div}\nolimits\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}-F({{\bf{d}}})\right)\cdot{{\bf{v}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot(-\Delta{{\bf{d}}}+f({{\bf{d}}}))\varphi$
(4.2) $\displaystyle=\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{I}}\underbrace{+\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot f({{\bf{d}}})\varphi}_{\mathcal{II}}.$
Differentiating (${}_{3}$ gives
(4.3) $\partial_{t}\nabla{{\bf{d}}}+\nabla\left[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}\right]=\Delta(\nabla{{\bf{d}}})-\nabla(f({{\bf{d}}})).$
Taking the inner product of (4.3) with respect to $\nabla{{\bf{d}}}\varphi$ and integrating over $\mathbb{T}^{3}$ we obtain
$\int_{\mathbb{T}^{3}}\partial_{t}(\nabla{{\bf{d}}}):\nabla{{\bf{d}}}\varphi=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\partial_{t}\varphi.$
Meanwhile, we have that
$\displaystyle\int_{\mathbb{T}^{3}}\nabla[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]:\nabla{{\bf{d}}}\varphi$ $\displaystyle=\int_{\mathbb{T}^{3}}-[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi-[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}$
$\displaystyle=-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}$
(4.4) $\displaystyle=\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{z}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{III}}\underbrace{-\int_{\mathbb{T}^{3}}[{{\bf{v}}}\cdot\nabla]{{\bf{d}}}\cdot\Delta{{\bf{d}}}\varphi}_{\mathcal{IV}}-\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{\bf{d}}}\cdot[\nabla\varphi\cdot\nabla]{{\bf{d}}}.$
Then through integration by parts we obtain for the right-hand side
(4.5) $\int_{\mathbb{T}^{3}}\nabla(\Delta{{\bf{d}}}-f({{\bf{d}}})):\nabla{{\bf{d}}}\varphi=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}|^{2}\varphi-\int_{\mathbb{T}^{3}}\Delta{{\bf{d}}}\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}})-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi$
$=$ $-\int_{\mathbb{T}^{3}}|\Delta{{\bf{d}}}|^{2}\varphi$
$+\int_{\mathbb{T}^{3}}\frac{1}{2}|\nabla{{\bf{d}}}|^{2}\Delta\varphi+\int_{\mathbb{T}^{3}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{{\bf{d}}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
$-\int_{\mathbb{T}^{3}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Finally, taking the inner product of (${}_{3}$ with $f({\bf{d}})\varphi$ yields
$\displaystyle\int_{\mathbb{T}^{3}}\partial_{t}{{\bf{d}}}\cdot f({{\bf{d}}})\varphi$ $\displaystyle=\frac{\,\mathrm{d}}{{\,\mathrm{d}t}}\int_{\mathbb{T}^{3}}F({{\bf{d}}})\varphi-\int_{\mathbb{T}^{3}}F({{\bf{d}}})\partial_{t}\varphi,$
(4.6) $\displaystyle\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot f({{\bf{d}}})\varphi$ $\displaystyle=\underbrace{\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla{\bf{d}})\cdot f({\bf{d}})\varphi}_{\mathcal{V}}\underbrace{+\int_{\mathbb{T}^{3}}({\bf{v}}\cdot\nabla{\bf{d}})\cdot f({\bf{d}})\varphi}_{\mathcal{VI}},$
(4.7) $\displaystyle\int_{\mathbb{T}^{3}}(\Delta{{\bf{d}}}-f({{\bf{d}}}))\cdot f({{\bf{d}}})\varphi$ $\displaystyle=-\int_{\mathbb{T}^{3}}\nabla{{\bf{d}}}:\nabla f({{\bf{d}}})\varphi+f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla{{\bf{d}}})+|f({{\bf{d}}})|^{2}\varphi.$
Combining all the above identities, we obtain (4.1). Particularly, we utilize the following cancellations
$\displaystyle\mathcal{I}-\mathcal{IV}=0,$
$\displaystyle\mathcal{II}-\mathcal{VI}=0.$
Moreover, note that we have
$\mathcal{III}+\mathcal{V}=-\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi$
which then yields the seventh term on the right-hand-side of (4.1) ∎
We note that the above computations still hold if we choose $\varphi$ to be a constant function and the same calculation will result in the following global energy equality:
Corollary 4.2.
Under the seme assumptions as in Lemma 4.1, it holds that
(4.8) $\frac{{\,\mathrm{d}}}{{\,\mathrm{d}}t}\int_{\mathbb{T}^{3}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{\mathbb{T}^{3}}\left(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}\right)$
$=\int_{\mathbb{T}^{3}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla]{{{\bf{v}}}}\cdot{{{\bf{z}}}}+\int_{\mathbb{T}^{3}}({\bf{z}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
4.2. Local-energy-preserving approximation
Given ${{\bf{z}}}(\omega)$ as in Lemma 3.2, we introduce the following approximated equations to (3.3):
(4.9) $\left\{\begin{array}[]{l}\partial_{t}{{\bf{v}}}^{\sigma}-\Delta{{\bf{v}}}^{\sigma}+({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla({{\bf{z}}}+{{\bf{v}}}^{\sigma})+\nabla{\pi}^{\sigma}=-\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})),\\ \nabla\cdot{{\bf{v}}}^{\sigma}=0,\\ \partial_{t}{{\bf{d}}}^{\sigma}-({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]=\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}),\end{array}\right.$
where $\sigma>0$ is the approximation parameter and $\Phi_{\sigma}[f]$ denotes the smooth mollification of a function $f$. More precisely, suppose that $\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$ is a standard mollifier supported in the $B_{1}(0)$, and let $\psi_{\sigma}:=\sigma^{-3}\psi(x/\sigma)$. Then we define $\Phi_{\sigma}[f]$ via the following convolution

### v2

@@ 3.2 Definitions of suitable weak solutions
Assume that $T>0$ and fix $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13). Then we say $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (11) in the sense of distribution on
$Q_{T}:=\mathbb{T}^{3}\times(0,T],$
subject to ${{\bf{v}}}(0)={{\bf{u}}}_{0},{{\bf{d}}}(0)={{\bf{d}}}_{0}$, if for any $\phi_{1}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, $\mathop{\rm div}\nolimits\phi_{1}=0$, $\phi_{2}\in C^{\infty}(\mathbb{T}^{3},\mathbb{R}^{3})$, and $t>0$, the following weak formulation of (11) is satisfied:
@@
Definition 2.
Assume that $T>0$ and fix $\omega\in\Omega$ such that ${\bf{z}}(\omega)$ satisfies the bound specified in (13). Then a suitable weak solution to (4) is a triple $({{\bf{u}}},P,{{\bf{d}}})$ such that if ${{\bf{v}}}={{\bf{u}}}-{{\bf{z}}}$ and ${\pi}=P-{Q}$ where $({{\bf{z}}},{Q})$ is the solution to (10), then
  1. 1.
@@
  3. 3.
    $({{\bf{v}}},{{\bf{d}}})$ is a weak solution to (11) in the sense of distribution on $Q_{T}$ as specified in Def. 1,
  4. 4.
    for all $t\leq T$ and almost every $s<t$, we have
    $\int_{\mathbb{T}^{3}\times\{t\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)+\int_{s}^{t}\int_{\mathbb{T}^{3}}|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}-f({{\bf{d}}})|^{2}$ (14)
    $\leq$ $\int_{\mathbb{T}^{3}\times\{s\}}\left(\frac{1}{2}|{{\bf{v}}}|^{2}+\frac{1}{2}|\nabla{{\bf{d}}}|^{2}+F({{\bf{d}}})\right)$
    $+\int_{s}^{t}\int_{\mathbb{T}^{3}}{{\bf{z}}}\cdot[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{v}}}]+({{\bf{z}}}\cdot\nabla){{\bf{d}}}\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}})).$
  5. 5.
    for any $\varphi\in C_{0}^{\infty}(Q_{T})$ (we note that the subscript “0” refers to compact support in the corresponding domain), $\varphi\geq 0$, we have
    $\int_{\mathbb{T}^{3}\times\{T\}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)+\int_{Q_{T}}(|\nabla{{\bf{v}}}|^{2}+|\Delta{{\bf{d}}}|^{2}+|f({{\bf{d}}})|^{2})\varphi$ (15)
    $\leq$ $\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}+F({{\bf{d}}})\right)\partial_{t}\varphi+\int_{Q_{T}}[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{\bf{v}}]\cdot{\bf{z}}\varphi$
    $+\int_{Q_{T}}\left(\frac{|{{\bf{v}}}|^{2}}{2}+{\bf{v}}\cdot{\bf{z}}\right)[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla\varphi]+{\pi}{{\bf{v}}}\cdot\nabla\varphi+\left(\frac{|{{\bf{v}}}|^{2}}{2}+\frac{|\nabla{{\bf{d}}}|^{2}}{2}\right)\Delta\varphi$
    $+\int_{Q_{T}}\left(\nabla{{\bf{d}}}\odot\nabla{{\bf{d}}}-\frac{1}{2}|\nabla{\bf{d}}|^{2}{\rm I}_{3}\right):\nabla^{2}\varphi$
    $+\int_{Q_{T}}({{\bf{z}}}\cdot\nabla{{\bf{d}}})\cdot(\Delta{{\bf{d}}}-f({{\bf{d}}}))\varphi+[({{\bf{z}}}+{{\bf{v}}})\cdot\nabla{{\bf{d}}}]\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}$
    $-\int_{Q_{T}}f({{\bf{d}}})\cdot(\nabla\varphi\cdot\nabla){{\bf{d}}}-2\int_{Q_{T}}\nabla f({{\bf{d}}}):\nabla{{\bf{d}}}\varphi.$
Definition 3.
[…]
$\widetilde{\mathfrak{B}}=(\widetilde{\Omega},\widetilde{\mathcal{F}},\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0},\widetilde{\mathbb{P}},(\widetilde{W}(t))_{t\geq 0}),$ (16)
and $(\widetilde{W}(t))_{t\geq 0}$ is a Brownian motion adapted to the filtration $\{\widetilde{\mathcal{F}}_{t}\}_{t\geq 0}$ with values in $D(A^{\delta})$ and covariance operator $\mathcal{L}$, such that
@@
is a $\widetilde{\mathcal{F}}_{t}$ adapted process, and
@@
is a measurable mapping, and there exists a set $\widetilde{\Omega}_{0}\subset\widetilde{\Omega}$ of full probability such that on this set the triple $)\displaystyle(\widetilde{{\bf{u}}}(\cdot,\widetilde{\omega}),\widetilde{P}(\cdot,\widetilde{\omega}),\widetilde{{\bf{d}}}(\cdot,\widetilde{\omega}$)) is a suitable weak solution in the sense of Definition 2.
3.3 Main results: existence and partial regularity
To state the main result of our article we introduce the notion of singular points for a suitable weak solution.
Definition 4.
@@ 3.3 Main results: existence and partial regularity
Our main result is as follows.
Theorem 2.
With the assumptions (9), suppose furthermore that the Brownian motion $\left(W(t)\right)_{t\geq 0}$ takes value in $D(A^{\frac{1}{4}+\delta})$ for some $\delta>0$. Then there exists a martingale suitable weak solution to (4) globally in time such that with full probability,
@@
4 Existence of martingale suitable weak solutions
@@
In this section, we will prove the first part of the conclusion of Theorem 2, that is, existence of martingale suitable weak solutions to (4) in the sense of Definition 3.
4.1 Energy-inequality-preserving approximation
We begin the proof of existence of martingale suitable weak solutions by introducing a system that converges to the SEL while preserving relevant energy inequalities.
Given ${{\bf{z}}}(\omega)$ as in Lemma 1, we introduce the following approximating equations to (11):
$\begin{cases}&\partial_{t}{{\bf{v}}}^{\sigma}-\Delta{{\bf{v}}}^{\sigma}+({{\bf{z}}}+\Phi_{\sigma}[{{\bf{v}}}^{\sigma}])\cdot\nabla({{\bf{z}}}+{{\bf{v}}}^{\sigma})+\nabla{\pi}^{\sigma}\\ &=-\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]\cdot(\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma})),\\ &\nabla\cdot{{\bf{v}}}^{\sigma}=0,\\ &\partial_{t}{{\bf{d}}}^{\sigma}-({{\bf{z}}}+{{\bf{v}}}^{\sigma})\cdot\nabla\Phi_{\sigma}[{{\bf{d}}}^{\sigma}]=\Delta{{\bf{d}}}^{\sigma}-f({{\bf{d}}}^{\sigma}),\end{cases}$ (17)
where $\sigma>0$ is the approximation parameter and $\Phi_{\sigma}[f]$ denotes the smooth mollification of a function $f$. More precisely, suppose that $\psi\in C_{0}^{\infty}(\mathbb{R}^{3})$ is a standard mollifier supported in $B_{1}(0)$, and let $\psi_{\sigma}:=\sigma^{-3}\psi(x/\sigma)$. Then we define $\Phi_{\sigma}[f]$ via the following convolution